# Research Paper

### v1

@@ 6.2 Accretion Rate & Line Width as Predictors of RM Lag
It is matter falling on to the accretion disk circling the SMBH that drives AGN activity, and there is evidence of a separation in behaviour between low and high accretion AGN. Highly accreting ‘super-Eddington’ AGN are associated with lower lags than their ‘sub-Eddington’ counterparts at the same luminosity (Du_2015). It is also evident that high accretion sources tend to sit below our best fit relation (e.g. Fig. 11), while low accretion sources sit above. We are then motivated to see if observations other than luminosity are predictors of AGN lag as a means to reduce the scatter of our single epoch methods, improve the predictive constraints of our single epoch methods and as a probe of the underlying AGN behaviour.
To examine this, we add emission line width as a variable in the relation, i.e. an relation:
@@
There is the possibility that the lack of predictive power comes from a strong correlation between luminosity and velocity dispersion in our sample: if we only probe a narrow band of space, the two parameters become degenerate in our fitting. We can rule this out by examining the correlation between luminosity and velocity (Fig. 13). We find that this correlation is weak, and is not consistent between lines. Table 6 notes an extremely shallow scaling index between luminosity and velocity dispersion for all lines, indicating that our sampling of the two parameters is sufficiently uncorrelated.
We note that this lack of predictive power in velocity is in contrast to the findings of the SEAMBH collaboration, (HBETA_Du_2016; HBETA_Du_2018; HBETA_Hu_2021) who find that higher accretion rates (derived from higher velocities) is associated with lower lags. Our results here do not necessarily contradict their findings: we do not include their data in these fits specifically because their highly accreting sources produce statistically lower H lags (see Section 6.1.1), nor do we use their critical threshold of accretion rate. Our fit only demonstrates that line width / accretion rate does not serve to explain the residuals about our best fits. It is also worth noting that different data sources have non-uniform approaches to measuring these line widths, and unlike the lag parameter we make no attempt to account for these differences in analysis.
| H | |||||||||
| MgII | |||||||||
| CIV | |||||||||
@@ 7 Probing BLR Stratification by Comparisons of Bolometric Relations
We use the bolometric corrections of Runnoe_2012, namely that , , and . We use the correction for for CIV, though note that the luminosity is defined at . As we take the logarithm of these values, small variations should have little impact.
@@
Performing this fit with luminosity units of (Fig. 18(a) (b)), we find that the MgII and H relations are very similar, though the MgII lags sit at a slightly higher offset and prefer a slightly shallower slope. By contrast, the fit for CIV sits lower, but with a slope and scatter that is broadly similar to H. This is in keeping with prior studies of sources with multiple lines visible simultaneously, which found that H and MgII were produced by roughly co-spatial regions of the BLR (i.e. yielded similar lags for the same source, for an example see MgII_Shen_2019), while CIV lags tend to be smaller by a factor of , though this ratio is poorly constrained. (MgII_Lira_2018; Kaspi_2007).
@@
Here, units for lag and luminosity are as in equation 6. This allows each line’s relation to have a different vertical offset, equivalent to the assumption that the ratio of these lags is the same across different AGN.
@@
Constraints on suggest MgII produces longer lags than H, indicating that it may be emitted from a radius larger than H by a factor of . This finding is statistically significant at . This is consistent with OzDES_Malik_2024 who did a stacked RM analysis (see their Table 1), in which they found that the average MgII lags in the OzDES sample were significantly longer than those for H at comparable luminosities.
@@
We find that the CIV radius is significantly smaller than that of H (i.e. ) by a factor of .
The constraints on the relation as well as the relative scaling factors are summarised in Table 7, with a sketch of the implications for the BLR geometry shown in Fig. 14.
| Data |
| H | - | - | |||||||||||||||||
| H + MgII | - | ||||||||||||||||||
| H + CIV | - | ||||||||||||||||||
| H + MgII + CIV | |||||||||||||||||||
@@ 8 Black Hole Masses From Single-Epoch Measurements (New Results)
Have a mass that obeys at a level
@@
As with the RM sources in Section 5.2, we also use equation 4 to estimate the accretion rates of these sources, using the bolometric corrections of Runnoe_2012 to convert monochromatic luminosities to the required units. All single epoch sources, including their masses, accretion rates, single epoch radii and measured velocity dispersions, are listed in Table LABEL:tab:single_epoch. These masses are included, along with the RM and literature measurements, in Fig. 4.
@@ 9 Discussion and Conclusions
Supplementing OzDES data with existing lags from the literature, we provide the best and most reliable constraints to date on the radius-luminosity relationship for sub-Eddington H and MgII lags. We are unable to fit a single relationship for CIV due to a disparity in the trends for low and high redshift sources. For all lines, we do not find a strong impact from accretion rate or emission line broadening as a single epoch lag predictor. Applying our MgII relationship to ULAS J1120+0641, a particularly distant SMBH observed by JWST at a redshift of (Banados_2017), we recover a mass of , slightly lower than Banados_2017’s estimate of .
@@
Our single epoch relations for H and MgII have a significantly smaller scatter than those of the final SDSS release, dex against their dex, offering a precise and broadly applicable tool for estimating SMBH masses to higher redshifts. We also offer the best constrained CIV fit at high redshift, but we recommend some caution in extrapolating this relation beyond the luminosity range of the available RM-data (i.e. ).
@@
We note that measurements we present are subject to a selection function that depends on interaction between the physical properties of the AGN, the varied analysis pipelines of different surveys and the quality cuts that we apply here. These effects are particularly strong at high redshift where lags abut our upper observational limit of observer-frame lags. We make no attempt to model these selection effects, and so defer any interrogation of the black hole population properties to future work.
We note that this ceiling on the AGN SMBH masses has lowered towards more recent times, with no SMBH above nearer than in our results. If we limit ourselves to only the DES single-epoch data of Table LABEL:tab:single_epoch we acquire a single homogeneous data set that avoids the observation baseline dependent effects discussed in Section 6.1.3. Even for this data set, we still see a significant evolution of the mass-distribution of sources as a function of redshift (Fig. 19), with nearby sources showing a consistent shift towards lower masses. As OzDES is a magnitude limited survey, some of the lack of distant low-mass observations can be explained by low completeness, dim objects are harder to find at greater distances, but we can still confirm that the high-mass tail of the mass-density distribution, which should be subject only to count-based statistical effects, also decreases for nearby times to a degree that cannot be explained by statistical uncertainty or by a decrease in the number of available sources.
@@
The similarity of the MgII and H relationship in slope and scatter, even with a statistically significant sample such as we now have access to, supports our understanding of their emission regions as being stratified. Of interest is the systematically larger MgII lags for comparable bolometric luminosity. Simple ionization models that form the basis for the relationship suggest the MgII emission line region should be located closer to the SMBH to the engine, rather than further. We note this with the caveat that our analysis is based on the simplest possible modelling of the spectral-energy distribution of AGN.
Conversely, our fits indicate that CIV is emitted from closer in than either H or MgII by a factor of , though the challenges in fitting a single relationship for this line make it difficult to constrain this with any certainty.
@@
The strong windowing effects of the CIV trend highlights the need for further study of this emission line,666For example, to investigate whether the CIV line is affected by outflows, which could contaminate the line’s velocity dispersion measurement. either through intensive study of low redshift sources or by extending the luminosity envelope of CIV lag recoveries. Nearby, low luminosity sources can be observed in the observer-frame ultraviolet range, (e.g. MgII_Metzroth_2006; MgII_Lira_2018), and extending such studies to higher redshift or luminosities would help to clarify the origins of the apparent flattening of the CIV curve. For the same reason, there is also motivation to use infrared spectroscopy to observe the high redshift counterparts to the H and MgII sample to determine if these systems are homogeneous or show evolution over time. Extending these samples to more distant, more luminous sources will allow us to determine whether this ‘levelling out’ of the relationship is physical, a statistical anomaly, or a selection effect. It will also let us probe the evolution of the geometry of such systems over time, and help delineate between accretion and luminosity based effects. In particular, follow up of sources in the low redshift domain, where the lower impact of time dilation means AGN variability is fast and lags are short, will allow direct comparison between CIV lags to H lags from and to MgII lags from . MgII_Metzroth_2006 and MgII_Lira_2018 already do this for some local AGN, but extending to more distant and luminous sources will help fill the low luminosity branch of the CIV relationship.
@@ Acknowledgements
Based in part on data acquired at the Anglo-Australian Telescope, under program A/2013B/012]. We acknowledge the traditional owners of the land on which the AAT stands, the Gamilaroi people, and pay our respects to elders past and present.
@@
1School of Mathematics and Physics, University of Queensland, Brisbane, QLD 4072, Australia, 2Kavli Institute for Particle Astrophysics & Cosmology, P. O. Box 2450, Stanford University, Stanford, CA 94305, USA, 3Center for Cosmology and Astro-Particle Physics, The Ohio State University, Columbus, OH 43210, USA, 4Department of Astronomy, The Ohio State University, Columbus, OH 43210, USA, 5Centre for Gravitational Astrophysics, College of Science, The Australian National University, ACT 2601, Australia, 6The Research School of Astronomy and Astrophysics, Australian National University, ACT 2601, Australia, 7Sydney Institute for Astronomy, School of Physics, A28, The University of Sydney, NSW 2006, Australia, 8School of Mathematical & Physical Sciences, 12 Wally’s Walk, Macquarie University, Macquarie Park, NSW 2113, 9Fermi National Accelerator Laboratory, P. O. Box 500, Batavia, IL 60510, USA, 10Physik-Institut, University of Zúrich, Winterthurerstrasse 190, CH-8057 Zúrich, Switzerland, 11Departamento de Física Teórica, Centro de Astropartículas y Física de Altas Energías (CAPA), Universidad de Zaragoza, 50009 Zaragoza, Spain, 12Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK, 13University Observatory, LMU Faculty of Physics, Scheinerstr. 1, 81679 Munich, Germany, 14Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK, 15Instituto de Astrofisica de Canarias, E-38205 La Laguna, Tenerife, Spain, 16Laboratório Interinstitucional de e-Astronomia - LIneA, Av. Pastor Martin Luther King Jr, 126 Del Castilho, Nova América Offices, Torre 3000/sala 817, Brazil, 17Universidad de La Laguna, Dpto. Astrofísica, E-38206 La Laguna, Tenerife, Spain, 18INAF-Osservatorio Astronomico di Trieste, via G. B. Tiepolo 11, I-34143 Trieste, Italy, 19Korea Astronomy and Space Science Institute, 776 Daedeok-daero, Yuseong-gu, Daejeon 34055, South Korea, 20Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona) Spain, 21Kapteyn Astronomical Institute, University of Groningen, Landleven 12 (Kapteynborg, 5419), 9747 AD Groningen, The Netherlands, 22Hamburger Sternwarte, Universität Hamburg, Gojenbergsweg 112, 21029 Hamburg, Germany, 23Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Madrid, Spain, 24California Institute of Technology, 1200 East California Blvd, MC 249-17, Pasadena, CA 91125, USA, 25Instituto de Física Teórica UAM/CSIC, Universidad Autónoma de Madrid, 28049 Madrid, Spain, 26Centre for Astrophysics & Supercomputing, Swinburne University of Technology, Victoria 3122, Australia, 27Santa Cruz Institute for Particle Physics, Santa Cruz, CA 95064, USA, 28Center for Astrophysics Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA, 29Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA, 30Lowell Observatory, 1400 Mars Hill Rd, Flagstaff, AZ 86001, USA, 31Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109, USA, 32Department of Physics and Astronomy, University of Pennsylvania, Philadelphia, PA 19104, USA, 33George P. and Cynthia Woods Mitchell Institute for Fundamental Physics and Astronomy, and Department of Physics and Astronomy, Texas A&M University, College Station, TX 77843, USA, 34Université Grenoble Alpes, CNRS, LPSC-IN2P3, 38000 Grenoble, France, 35Center for Astrophysical Surveys, National Center for Supercomputing Applications, 1205 West Clark St., Urbana, IL 61801, USA, 36Department of Astronomy, University of Illinois at Urbana-Champaign, 1002 W. Green Street, Urbana, IL 61801, USA, 37Institució Catalana de Recerca i Estudis Avançats, E-08010 Barcelona, Spain, 38Department of Astrophysical Sciences, Princeton University, Peyton Hall, Princeton, NJ 08544, USA, 39Observatório Nacional, Rua Gal. José Cristino 77, Rio de Janeiro, RJ - 20921-400, Brazil, 40Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute, German Centre for Cosmological Lensing, 44780 Bochum, Germany, 41Physics Department, Lancaster University, Lancaster, LA1 4YB, UK, 42Computer Science and Mathematics Division, Oak Ridge National Laboratory, Oak Ridge, TN 37831, 43Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL 60637, USA, 44Cerro Tololo Inter-American Observatory, NSF’s National Optical-Infrared Astronomy Research Laboratory, Casilla 603, La Serena, Chile, 45Berkeley Center for Cosmological Physics, Department of Physics, University of California, Berkeley, CA 94720, US

### v2

@@ 6.2 Accretion Rate & Line Width as Predictors of RM Lag
It is matter falling on to the accretion disk circling the SMBH that drives AGN activity, and there is evidence of a separation in behaviour between low and high accretion AGN. Highly accreting ‘super-Eddington’ AGN are associated with lower lags than their ‘sub-Eddington’ counterparts at the same luminosity (Du_2015). It is also evident that high accretion sources tend to sit below our best fit relation (e.g. Figure 12), while low accretion sources sit above. We are then motivated to see if observations other than luminosity are predictors of AGN lag as a means to reduce the scatter of our single epoch methods, improve the predictive constraints of our single epoch methods and as a probe of the underlying AGN behaviour.
To examine this, we examine a new functional form for the single epoch lag prediction in which we add emission line width as a variable in the relation, i.e. an relation:
@@
There is the possibility that the lack of predictive power comes from a strong correlation between luminosity and velocity dispersion in our sample: if we only probe a narrow band of space, the two parameters become degenerate in our fitting. We can rule this out by examining the correlation between luminosity and velocity (Figure 14). We find that this correlation is weak, and is not consistent between lines. Table 6 notes an extremely shallow scaling index between luminosity and velocity dispersion for all lines, indicating that our sampling of the two parameters is sufficiently uncorrelated.
We note that this lack of predictive power in velocity is in contrast to the findings of the SEAMBH collaboration, (HBETA_Du_2016; HBETA_Du_2018; HBETA_Hu_2021) who find that higher accretion rates (derived from higher velocities) is associated with lower lags. Our results here do not necessarily contradict their findings, this analysis is for our primary data group which does not include their data specifically because their highly accreting sources produce statistically lower H lags (see Section 6.1.1), and we use a different functional form for the accretion dependence in place of their use of a critical accretion rate threshold. Our fit only demonstrates that line width / accretion rate does not serve to explain the residuals about our best fits. It is also worth noting that different data sources have non-uniform approaches to measuring these line widths, and unlike the lag parameter we make no attempt to account for these differences in analysis.
| Line Type | Log-Vel / Log-Lum Slope | Scatter (Dex) |
| H | ||
| MgII | ||
| CIV | ||
@@ 7 Probing BLR Stratification by Comparisons of Bolometric Relations
We use the bolometric corrections of Runnoe_2012, namely that , , and . We use the correction for for CIV, though note that the luminosity is defined at . As we take the logarithm of these values, small variations should have little impact. The resulting bolometric luminosities span the range .
@@
Performing this fit with luminosity units of (Figure 19(a) (b)), we find that the MgII and H relations are very similar, though the MgII lags sit at a slightly higher offset and prefer a slightly shallower slope. By contrast, the fit for CIV sits lower, but with a slope and scatter that is broadly similar to H. This is in keeping with prior studies of sources with multiple lines visible simultaneously, which found that H and MgII were produced by roughly co-spatial regions of the BLR (i.e. yielded similar lags for the same source, for an example see MgII_Shen_2019), while CIV lags tend to be smaller by a factor of , though this ratio is poorly constrained. (MgII_Lira_2018; Kaspi_2007).
@@
Here, units for lag and luminosity are as in equation 6. This allows each line’s relation to have a different vertical offset, equivalent to the assumption that the ratio of these lags is the same across different AGN. 666A further model was trialled in which each line was free to fit its own inherent scatter, but this did not significantly change the resulting parameter constraints for slope or scale factors compared to when a single intrinsic scatter is fit.
@@
Constraints on suggest MgII produces longer lags than H, indicating that it may be emitted from a radius larger than H by a factor of . OzDES_Malik_2024 who did a stacked RM analysis (see their Table 1), in which they found that the average MgII lags in the OzDES sample were significantly longer than those for H at comparable luminosities.
@@
We find that the CIV radius is consistent with begin co-spatial with the H radius, though it prefers to be smaller by a factor of
The constraints on the relation as well as the relative scaling factors are summarised in Table 7, with a sketch of the implications for the BLR geometry shown in Figure 15. These findings are in line with those of SDSS-Shen_2023, who similarly found MgII lags to be systematically longer than those for H and CIV lags to be systematically shorter.
| Data | Slope () | Offset () | Scatter () | Slope - Offset Fit Correlation () | ||
| H | - | - | ||||
| H + MgII | - | |||||
| H + CIV | - | |||||
| H + MgII + CIV | ||||||
@@ 8 Black Hole Masses From Single-Epoch Measurements (New Results)
Have a mass that excludes a nonphysical zero/negative mass at a level
@@
As with the RM sources in Section 5.2, we also use equation 4 to estimate the accretion rates of these sources, using the bolometric corrections of Runnoe_2012 to convert monochromatic luminosities to the required units. All single epoch sources, including their masses, accretion rates, single epoch radii and measured velocity dispersions, are listed in Table LABEL:tab:single_epoch. These masses are included, along with the RM and literature measurements, in Figure 5.
@@
We find that the ceiling on the AGN SMBH masses has lowered towards more recent times, with no SMBH above nearer than in our results. If we limit ourselves to only the DES single-epoch data of Table LABEL:tab:single_epoch we acquire a single homogeneous data set that avoids the observation baseline dependent effects discussed in Section 6.1.3. Even for this data set, we still see a significant evolution of the mass-distribution of sources as a function of redshift (Figure 19), with nearby sources showing a consistent shift towards lower masses. As OzDES is a magnitude limited survey, some of the lack of distant low-mass observations can be explained by low completeness, dim objects being harder to observe at greater distances, but we can still confirm that the high-mass tail of the mass-density distribution, which should be subject only to count-based statistical effects, also decreases for nearby times to a degree that cannot be explained by statistical uncertainty or by a decrease in the number of available sources (see Figure 20). This is not a shocking result, as it aligns with the well established trend of ‘cosmic downsizing’ by which massive SMBH become less active towards more recent cosmic history (Barger_2005_downsizing; Vestergaard_2009_downsizing; Kelly_2010_downsizing; Fanidakis_2011_downsizing).
@@ 9 Discussion and Conclusions
Supplementing OzDES data with existing lags from the literature that are consistent OzDES’s high-redshift sources, we provide significantly improved constraints on the radius-luminosity relationship for sub-Eddington H and MgII lags. For all lines, we do not find a strong impact from accretion rate or emission line broadening as a single epoch lag predictor. Applying our MgII relationship to ULAS J1120+0641, a particularly distant SMBH observed by JWST at a redshift of and a luminosity of and MgII line FWHM of (Banados_2017), we recover a mass of , slightly lower, but still consistent with, Banados_2017’s estimate of .
@@
Our single epoch relations for H and MgII have a significantly smaller scatter than those of the final SDSS release, dex against their dex. As these relations are tuned over a large redshift range, they offer a broadly applicable tool for estimating SMBH masses to higher redshifts. We also offer the best constrained CIV fit at high redshift, but we recommend some caution in extrapolating this relation beyond the luminosity range of the available RM-data (i.e. ).
@@
Using our new relations, we make single epoch mass estimates for AGN up to redshift . In these estimates we are able to observe a dearth of massive AGN, with the upper range of AGN mass trending downwards at recent cosmic times. We note that measurements we present are subject to a selection function that depends on interaction between the physical properties of the AGN, the varied analysis pipelines of different surveys and the quality cuts that we apply here. These effects are particularly strong at high redshift where lags abut our upper observational limit of observer-frame lags. We make no attempt to model these selection effects, and so defer any interrogation of the black hole population properties to future work.
@@
The similarity of the MgII and H relationship in slope and scatter, even with a statistically significant sample such as we now have access to, supports our understanding of their emission regions as being stratified. Of interest is the systematically larger MgII lags for comparable bolometric luminosity. Though simple arguments of ionisation energy suggest that the higher energy MgII should precipitate out closer in than H, some photoionisation models (e.g. Guo_2020_Photoionization) do support MgII being farther out from the SMBH engine. We note this with the caveat that our analysis is based on the simplest possible modelling of the spectral-energy distribution of AGN. Conversely, our fits indicate that CIV may be emitted from a radius much closer to that of H than found from simultaneous RM of H and CIV (MgII_Lira_2018; Kaspi_2007), though a slightly smaller radius is statistically preferred. The challenges in fitting a single relationship for this line make it difficult to constrain this with any certainty.
@@
The strong windowing effects of the CIV trend highlights the need for further study of this emission line,777For example, to investigate whether the CIV line is affected by outflows, which could contaminate the line’s velocity dispersion measurement. either through intensive study of low redshift sources or by extending the luminosity envelope of CIV lag recoveries. Nearby, low luminosity sources can be observed in the observer-frame ultraviolet range, (e.g. MgII_Metzroth_2006; MgII_Lira_2018), and extending such studies to higher redshift or luminosities would help to clarify the origins of the apparent flattening of the CIV curve. For the same reason, there is also motivation to use infrared spectroscopy to observe the high redshift counterparts to the H and MgII sample to determine if these systems are homogeneous or show evolution over time. Extending these samples to more distant, more luminous sources will allow us to determine whether this ‘levelling out’ of the relationship is physical, a statistical anomaly, or a selection effect. It will also let us probe the evolution of the geometry of such systems over time, and help delineate between accretion and luminosity based effects. In particular, follow up of sources in the low redshift domain, where the lower impact of time dilation means AGN variability is fast and lags are short, will allow direct comparison between CIV lags to H lags from and to MgII lags from . MgII_Metzroth_2006 and MgII_Lira_2018 already do this for some local AGN, but extending to more distant and luminous sources will help fill the low luminosity branch of the CIV relationship.
The major findings of this paper are twofold: firstly that we have enough data as of this generation in BLR RM to make meaningful scaling relationships for H, MgII and CIV out to high redshift, and that the large apparent disagreement between surveys can be resolved with even a simple accounting of selection effects and a consistent framework for screening false positives. In the resulting relations the inherent scatter now vastly outstrips statistical uncertainty, meaning the next generation of BLR RM may concern itself with the second order effects of sub-populations, multi-variable lag predictors (e.g. accretion rate), sub-populations and more granular statistical biases. The second finding, and one we make an effort to stress, is that the handling of these selection effects and false positives are not a fully settled issue.
Though BLR RM surveys are conceptually simple, surveys can adopt wildly divergent approaches in the particulars of their lag recovery. Different regimes for fitting lags and different criteria for quantifying the significance of these fits interact with initial survey target selection in a complex and multi-layered way. In this paper we offer relations that are a significant improvement over the existing standard, specifically for high redshift sources. However there remains work to be done in future to improve these constraints, even with existing RM measurements. There are two main avenues of interest for strengthening our RM-derived scaling relations in their current form: firstly, to re-fit lags for all sources across all surveys in a single framework, and secondly to properly model how this framework interacts with surveys to understand what selection functions arise in the luminosity / redshift / lag plane. We suggest that characterising these window functions should include a combination of simulation-based forward modelling in the style of (OzDES-Penton_2021) with modern fitters like LITMUS and PyROA do not suffer from JAVELIN’s numerical artefacts when encountering aliasing (see McDougall_2025_LITMUS). Of particular interest to the OzDES data would be a re-analysis that discards less false negatives for a larger final sample, and understanding how such cuts affects the apparent population distributions in addition to the overall false positive rate.
@@ Acknowledgements
Based in part on data acquired at the Anglo-Australian Telescope, under program Ab/2013B/012]. We acknowledge the traditional owners of the land on which the AAT stands, the Gamilaroi people, and pay our respects to elders past and present.
@@
1School of Mathematics and Physics, University of Queensland, Brisbane, QLD 4072, Australia, 2Kavli Institute for Particle Astrophysics & Cosmology, P. O. Box 2450, Stanford University, Stanford, CA 94305, USA, 3Center for Cosmology and Astro-Particle Physics, The Ohio State University, Columbus, OH 43210, USA, 4Department of Astronomy, The Ohio State University, Columbus, OH 43210, USA, 5Centre for Gravitational Astrophysics, College of Science, The Australian National University, ACT 2601, Australia, 6The Research School of Astronomy and Astrophysics, Australian National University, ACT 2601, Australia, 7Sydney Institute for Astronomy, School of Physics, A28, The University of Sydney, NSW 2006, Australia, 8School of Mathematical & Physical Sciences, 12 Wally’s Walk, Macquarie University, Macquarie Park, NSW 2113, 9Fermi National Accelerator Laboratory, P. O. Box 500, Batavia, IL 60510, USA, 10Physik-Institut, University of Zúrich, Winterthurerstrasse 190, CH-8057 Zúrich, Switzerland, 11Departamento de Física Teórica, Centro de Astropartículas y Física de Altas Energías (CAPA), Universidad de Zaragoza, 50009 Zaragoza, Spain, 12Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK, 13University Observatory, LMU Faculty of Physics, Scheinerstr. 1, 81679 Munich, Germany, 14Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK, 15Instituto de Astrofisica de Canarias, E-38205 La Laguna, Tenerife, Spain, 16Laboratório Interinstitucional de e-Astronomia - LIneA, Av. Pastor Martin Luther King Jr, 126 Del Castilho, Nova América Offices, Torre 3000/sala 817, Brazil, 17Universidad de La Laguna, Dpto. Astrofísica, E-38206 La Laguna, Tenerife, Spain, 18INAF-Osservatorio Astronomico di Trieste, via G. B. Tiepolo 11, I-34143 Trieste, Italy, 19Korea Astronomy and Space Science Institute, 776 Daedeok-daero, Yuseong-gu, Daejeon 34055, South Korea, 20Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona) Spain, 21Kapteyn Astronomical Institute, University of Groningen, Landleven 12 (Kapteynborg, 5419), 9747 AD Groningen, The Netherlands, 22Hamburger Sternwarte, Universität Hamburg, Gojenbergsweg 112, 21029 Hamburg, Germany, 23Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Madrid, Spain, 24California Institute of Technology, 1200 East California Blvd, MC 249-17, Pasadena, CA 91125, USA, 25Instituto de Física Teórica UAM/CSIC, Universidad Autónoma de Madrid, 28049 Madrid, Spain, 26Centre for Astrophysics & Supercomputing, Swinburne University of Technology, Victoria 3122, Australia, 27Santa Cruz Institute for Particle Physics, Santa Cruz, CA 95064, USA, 28Center for Astrophysics Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA, 29Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA, 30Lowell Observatory, 1400 Mars Hill Rd, Flagstaff, AZ 86001, USA, 31Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109, USA, 32Department of Physics and Astronomy, University of Pennsylvania, Philadelphia, PA 19104, USA, 33George P. and Cynthia Woods Mitchell Institute for Fundamental Physics and Astronomy, and Department of Physics and Astronomy, Texas A&M University, College Station, TX 77843, USA, 34Université Grenoble Alpes, CNRS, LPSC-IN2P3, 38000 Grenoble, France, 35Center for Astrophysical Surveys, National Center for Supercomputing Applications, 1205 West Clark St., Urbana, IL 61801, USA, 36Department of Astronomy, University of Illinois at Urbana-Champaign, 1002 W. Green Street, Urbana, IL 61801, USA, 37Institució Catalana de Recerca i Estudis Avançats, E-08010 Barcelona, Spain, 38Department of Astrophysical Sciences, Princeton University, Peyton Hall, Princeton, NJ 08544, USA, 39School of Mathematics and Physics, University of Surrey, Guildford, Surrey, GU2 7XH, UK, 40Observatório Nacional, Rua Gal. José Cristino 77, Rio de Janeiro, RJ - 20921-400, Brazil, 41Ruhr University Bochum, Faculty of Physics and Astronomy, Astronomical Institute, German Centre for Cosmological Lensing, 44780 Bochum, Germany, 42Physics Department, Lancaster University, Lancaster, LA1 4YB, UK, 43Computer Science and Mathematics Division, Oak Ridge National Laboratory, Oak Ridge, TN 37831, 44Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL 60637, USA, 45Cerro Tololo Inter-American Observatory, NSF’s National Optical-Infrared Astronomy Research Laboratory, Casilla 603, La Serena, Chile, 46Berkeley Center for Cosmological Physics, Department of Physics, University of California, Berkeley, CA 94720, US